Calculate the exact volume of a square prism if the diagonal of the square face is 5 cm and the height is 2sqr2.
Thanks

Answer by CubeyThePenguin(3113)

side length = 5/sqrt(2) cm = 5sqrt(2)/2 cm

volume = (5sqrt(2)/2)^2 * (2sqrt(2)) = 25/2 * 2sqrt(2) = 25sqrt(2) cm^3